OpenStudy (anonymous):

5/8x + 1/16x=3/16 + x

OpenStudy (lilg132):

5/8 + 1/16 0.625 + 0.0625 = 0.6875 0.6875x = 3/16 + x 0.6875x - x = -0.3125 -0.3125x = 3/16 -0.3125x = 0.1875 x = -0.6 dont forget to click good answer :)

OpenStudy (anonymous):

- lol, thanks love !